A thin glass of refractive index 1.5 has optical power -5D in air . What will be the optical power of the same lens immersed in a liquid of refractive index 1.6 

Grade:10

1 Answers

Aditya kashyap
6 years ago
P= 1/f = (U-1)(1/r1 - 1/r2) is lens maker formula .... where U is refractive index of medium.. Now p1/p2= (Uga-1)(1/r1 - 1/r2) /(Ugl-1)(1/r1 - 1/r2) ... As r1 and r2 is same in both medium...So p1/p2= (Uga-1)/(Ugw-1) .... p1/p2= (Ug/Ua-1)/(Ug/Uw-1)...Here, p1 =-5,Ug=1.5,Ua=1,Uw=1.6.... -5/P2=(1.5 -1)/(1.5/1.6 -1).. -5/P2=(0.5)/(-0.1/1.6).... P2= 1/1.6 ..... P2 =0.625 D.